OpenStudy (anonymous):

Determine the values of 'a' for which the system has no solutions, exactly one solution, or infinitely many solutions: x + 2y - 3z = 4 3x - y + 5z = 2 4x + y + (a^2 - 14)z = a+2

OpenStudy (anonymous):

\[x + 2y - 3z = 4\]\[3x - y + 5z = 2\]\[4x + y + (a^2-14)z = a+2\]

OpenStudy (cggurumanjunath):